OpenStudy (anonymous):

In 1991, the life expectancy of males in certain a country was 63.5 years. In 1995, it was 67 years. Let E represent the life expectancy in years t and let t represent the number of years since 1991. The linear function E(t) that fits data is E(t)=?t +? Round to the nearest tenth

OpenStudy (anonymous):

first try to find the slope of the line. from 1991 to 1995 life exp. has raised by 3.5 years. meaning that x changed 4 meanwhile y changed 3.5 so the slope is 3.5/4=7/8 at 1991 t=0 but the function has to give 63.5 So the line has to be E(t)=63.5+(7/8)t